49.7 = (2x2)/((1-x)(2-x)) turns into 0.920x2-3.00x+2.00=0

Please show how this equation turned into that one.

Thanks
Answer #1

solution viven: 49.7- 222 (1-2)(2-2) 49.7 = 22² (0²-32+2) 49.7 (2²3n+2) = 22² → divided both side by 49.7 49.7 (2²-3n+2) = 22